Project

General

Profile

Actions

Ops/Comm Scheduling meeting

Location : CR11 R3 @ 15:30 pm on Tuesday 23rd

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Agenda

  • Actions from last meeting
    • Pierrick/Mark: Ensure that the pressure and material in the absorber volumes are recorded
    • Mark: Make more accurate measurements of tracker patch panel helium leak
    • Pierrick: Progress on monitoring O2 content of tracker with sensors. Monitoring He levels at TOFs?
    • John G: Resin test progress? (Re: tracker helium leak)
    • Durga/John N: What is the effect on the beamline of removing GVA1?
    • Pierrick: Feedback to alarm handler re: corrupted data was showing false positives
    • Pierrick: Shared services (e.g. water, air) to be included in alarm handler
    • Durga: Convergence on data quality flags needed
    • Durga: Progress on testing of reconstruction quality flags
  • DC report - activities in the Hall this week
  • MOM report - Data-taking this week
    • Lessons learnt during startup
    • Data plots from recent shifts
    • Summary of runs (MOM)
  • Decay Solenoid PS return and calibration
    • Kaeser compressor status
  • Focus coil and absorber installation
  • Controls and monitoring
    • Beamstop summary shows beamstop as closed when it's open
    • Radiation monitor EPICS display freezes for ~2min
    • Proton absorber switch readout doesn't update
    • Q5 and Q6 temperature readback still swapped
  • Scheduling of required hall work and operations
  • List of things to address by February run
  • AOB
    • Surveying of TOF1/TOF2/KL/EMR
    • Tracker waveguide improvements
  • Chair of meetings for next 5 weeks
    • Steve is teaching at this time for the next 4 weeks. DC will chair the meeting.
  • DONM : 1/3/2016

Minutes

Present: Paolo, Melissa, Andy N., Paul K., Mike C., Henry, John G., Scott W., Colin, Roy, Alan G., Paul S., Jaroslaw, Yordan, Chris R., Ryan, Ed, Pierrick, Vicky B., Josef, Victoria

  • O2 monitoring on SSD is done (0.5% oxygen in downstream tracker). Monitor sits before the bubbler.
    • Plan to take data with one tracker in air and one in helium (Action: Melissa/Ed)
    • SSU O2 monitoring needs connecting (Action: Craig)
  • Decision taken to install FC in beamline and evacuate the beampipe on Wed/Thursday
  • Decay Solenoid
    • PSU not yet at customs, arrival prob. next week
    • Mike & Vicky will think about warming up the cold box to check concerns about the compressor
  • Controls & Monitoring
    • Action: Pierrick/Paolo, consider increasing the delay before starting EPICS IOC's after a reboot
    • Q5 and Q6 temperature monitoring are now the correct way around
    • Action: Perrick, investigate what's wrong with the neutron monitoring IOC (reading pauses for several seconds, then quickly catches up, then pauses, etc.)
  • Hall scheduling
    • ISIS need access to turn off a chiller tomorrow morning (John G. organising)
    • TOF1, TOF2/KL/EMR survey to be arranged (Action: John G)
    • Beam protection system tests need scheduling (Action: Craig to liase with Mark Arnold)
    • Camhall03's picture is upside down
  • AOB
    • Anyone with an O2 phone should give the number to John G. It can then be added to the signal booster in the hall.
    • Missing data sections in the archiver (Action: Paolo to talk to Mike C./Vicky B. and find out what's missing and why)

Updated by Blackmore, Victoria almost 8 years ago ยท 13 revisions